11) When discussing bonds, convexity relates to the ________.

A. shape of the bond price curve
B. shape of the yield curve
C. slope of the yield curve
D. shape of the bond dealer

12) A zero-coupon bond has a yield to maturity of 5% and a par value of $1,000. If the bond matures in 16 years, it should sell for a price of __________ today.

A. $458.00
B. $641.00
C. $789.00
D. $1,100.00

13) The yield-to-maturity (YTM) on a bond is exactly what the investor will earn when buying and holding the bond to maturity:

A. if there is no default on the bond, and coupons are re-invested back into the market at the YTM

B. if there yield curve of rates remains upward-sloping instead of becoming inverted

C. if interest rates in the economy decline steadily while the investor is holding the bond.

D. none of the above

14) An investor is likely to exercise his/her rights on a convertible bond if:

A. the YTM on the bond increases

B. the stock price of the bond issuer increases significantly

C. inflation increases in the economy

D. there is default on the bond.

15) A zero-coupon bond:

A. typically trades at a discount to face value

B. would trade at a premium to face value in the rare circumstances that the bond has a negative yield

C. both (A) and (B) are true.

D. none of the above
Answer #1

11) C is correct.

12) PV = FV / (1 + r)^n = 1000 / 1.05^16 = $458

13) A is correct.

14) B is correct as the value of equity will be higher than the value of the bond.

15) C is correct. PV = FV / (1 + r)^n is the formula for zero coupon bond.